Two planes leave Chicago - one traveling due east and the other due west. The first travels 100 miles per hour faster than the second. How long will it be until they are 2000 miles apart?
Answer by Alan3354(21605)

Two planes leave Chicago - one traveling due east and the other due west. The first travels 100 miles per hour faster than the second. How long will it be until they are 2000 miles apart?
----------------
v = slower plane's speed
v+100 = faster plane's speed
Since they're going opposite directions, their relative speed (with respect to each other) is the sum, 2v+100
-------------
d = rt
2000 = (2v+100)*t
t = 2000/(2v+100) hours
t = 1000/(v+50) hours
That's all that can be done without additional info.